### v1

# FoundationGrasp: Generalizable Task-Oriented Grasping with Foundation Models

Abstract
Task-oriented grasping (TOG), which refers to the problem of synthesizing grasps on an object that are configurationally compatible with the downstream manipulation task, is the first milestone towards tool manipulation. Analogous to the activation of two brain regions responsible for semantic and geometric reasoning during cognitive processes, modeling the complex relationship between objects, tasks, and grasps requires rich prior knowledge about objects and tasks. Existing methods typically limit the prior knowledge to a closed-set scope and cannot support the generalization to novel objects and tasks out of the training set. To address such a limitation, we propose FoundationGrasp, a foundation model-based TOG framework that leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Comprehensive experiments are conducted on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset, demonstrating the superiority of FoudationGrasp over existing methods when generalizing to novel object instances, object classes, and tasks out of the training set. Furthermore, the effectiveness of FoudationGrasp is validated in real-robot grasping and manipulation experiments on a 7 DoF robotic arm. Our code, data, appendix, and video are publicly available at https://sites.google.com/view/foundationgrasp.
Note to Practitioners
This research is motivated by the problem of generalizable task-oriented grasping skill learning. Solving such a problem would significantly improve the robot’s level of automation and intelligence in tool manipulation for household and industrial tasks. Existing methods cannot handle unseen objects and tasks in open-world environments. To address such a problem, we propose to leverage the open-ended knowledge from foundation models to improve the generalization capability of existing TOG methods. This way, the robot can perform TOG w.r.t. unseen objects and tasks, facilitating downstream tool manipulation. The proposed method currently follows a sampling-and-evaluation pipeline. Such a design makes it highly dependent on the quality of the off-the-shelf grasp sampler. Future work aims to construct a one-stage TOG pipeline with improved robustness and efficiency. Generally, this research can be applied to diverse household and industrial scenarios involving tool manipulation, such as kitchenware cleaning and part assembly.
Index Terms:
Robotic Grasping, Perception for Grasping and Manipulation, Deep Learning in Grasping and Manipulation
I Introduction
Endowing intelligent robots with the ability of tool manipulation, close to the human level, opens up the potential for robots to accomplish diverse household and industrial tasks, such as part assembly and kitchenware cleaning. As an initial step towards tool manipulation, task-oriented grasping (TOG) necessitates the robot synthesizing grasps that are configurationally compatible with downstream manipulation tasks. Task-incompatible grasping will inevitably limit the success of manipulation tasks or, in some cases, pose physical risks to the user (e.g., hand over a knife by grasping the handle and exposing the sharp blade to the receiver).
Despite the significance of TOG, such skills are not yet readily accessible to robots nowadays. Considering a household robot operating in environments, such as apartments and garages, it needs to handle an open set of elements, including objects (instances, classes) and tasks, most of which are novel elements unseen during training. A naive solution is to collect task-oriented grasping data for each object with respect to each task in a brute-force manner. However, such a solution is infeasible in terms of manpower and time costs in open-world environments. In practical terms, one would like to train the robot only on a limited set of objects and tasks, while the robot is able to generalize the learned TOG skills to novel elements beyond the training examples.
In cognitive psychology, it is observed that two brain regions in the prefrontal cortex, which are respectively responsible for semantic and geometric reasoning, are jointly activated during cognitive processes[1, 2]. Inspired by this observation, recent works [3, 4, 5, 6, 7] have proposed incorporating semantic and geometric knowledge into TOG pipelines to improve the robots’ adaptability to different situations. For example, Murali et al.[5] contribute the largest and most diverse TOG dataset, TaskGrasp, and build a knowledge graph (KG) capturing semantic relationships between objects and tasks within the TaskGrasp dataset. In parallel, a series of works [6, 7] consider both semantic and geometric properties of pre-defined sets of objects and tasks. While the aforementioned methods have exhibited a degree of success, they experience significant performance degradation when handling novel elements out of the pre-defined sets. This can be naturally attributed to the fact that the prior knowledge is limited to a closed-set scope. While such a limitation may be acceptable under the closed-world assumption, a household robot often operates in open-world environments, thus resulting in unsatisfying TOG performance in real-world applications.
The above discussions have led us to the question: How can we endow robots with open-ended knowledge about objects and tasks to learn generalizable TOG skills? Recent advancements in foundation models (e.g., Large Language Model (LLM) and Vision-Language Model (VLM)) have enabled robots to seamlessly extract and harness the open-ended knowledge embedded in these models for physically grounded tasks [8, 9, 10, 11, 12]. Motivated by this trend, we propose FoundationGrasp, a foundation model-based TOG framework. As is depicted in Figure LABEL:fig:concept, in contrast to previous works that restrict the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills.
The pipeline of the proposed FoundationGrasp is as follows. When presented with an object (e.g., mug), a task (e.g., pour), and a language instruction (e.g., pour water from the mug), where the object and the task can be novel elements out of the training set, FoundationGrasp first prompts an LLM to generate semantic and geometric descriptions of the object and the task. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their feature representations. In the final stage, we draw inspiration from cognitive psychology and introduce a Transformer-based[13] task-oriented grasp evaluator with semantic and geometric branches. Evaluation on the contributed Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset demonstrates that FoundationGrasp outperforms existing TOG methods when generalizing to novel object instances, object classes, and tasks out of the training set. To validate the effectiveness in real-world applications, we further deploy FoundationGrasp on a Kinova Gen3 7 DoF robotic arm for task-oriented grasping and manipulation.
This study extends a preliminary version GraspGPT [14], which showcases the potential of improving the generalization capability of TOG with an LLM in a proof-of-concept manner. Specifically, GraspGPT focuses on the semantic aspect of TOG and leverages the transferability of language to generalize to novel object classes and tasks. Building on top of GraspGPT, this extended version follows the same idea and introduces an enhanced framework, FoundationGrasp, which leverages both semantic and geometric knowledge from foundation models to learn generalizable TOG skills. More importantly, we include additional baselines, experiments, and analysis to thoroughly demonstrate the value of combining foundation models in solving TOG problems and provide insights into the limitations. An extended discussion of related work on task-agnostic grasping, task-oriented grasping, and the applications of foundation models in robotics is also presented.
The main contributions of this paper are as follows:
  • We propose FoundationGrasp, a foundation model-based TOG framework that leverages the open-ended knowledge from foundation models to learn generalizable TOG skills.
  • To support the training and evaluation of FoundationGrasp, we contribute the Language and Vision Augmented TaskGrasp (LaViA-TaskGrasp) dataset.
  • A physical system is built to enable the robot to perform task-oriented grasping with novel objects and tasks guided by users’ natural language instructions.
The rest of this paper is organized as follows. In Section II, we review related work. Section III provides a problem formulation, and Section IV describes the details of the proposed FoundationGrasp framework. In Section V and Section VI, the experimental setup, experiment results, and analysis are presented. Finally, we discuss the limitations in Section VII and conclude the paper in Section VIII.
II Related Work
II-A Task-Agnostic Grasping
Robotic grasping has been a fundamental problem in robotics. Earlier studies optimize grasp quality to output analytic solutions satisfying stability constraints, such as force closure and form closure [15]. However, a significant drawback of the analytic approach is the need for complete object and hand models [16], hindering the generalization to novel objects.
With the rise of machine learning techniques in recent years, the field of grasp synthesis and planning has been revolutionized by data-driven methods [17, 18, 19, 20, 21] trained on various types of sensory data, including RGB-D images, depth images, and point clouds. For example, Dex-Net 2.0[17] and DeepGrasp[18] employ convolutional neural networks (CNN) to process RGB-D images and predict SE(2) planar grasp poses. PS-CNN [20] takes advantage of both analytic and data-driven approaches by performing primitive shape decomposition on depth images to generate 6 DoF grasp poses in SE(3) space. [19, 21] approach the problem of grasp synthesis on point clouds with sampling-and-evaluation pipelines. While a substantial portion of research in robotic grasping focuses on task-agnostic grasping (TAG), the goal of grasping is often beyond this realm. Ignoring the task-oriented effect on object grasping limits the application of TAG methods in many cases. FoundationGrasp complements TAG methods by predicting grasps satisfying both stability and task-compatibility constraints.
II-B Task-Oriented Grasping
In terms of TOG, where the grasping problem is contextualized in manipulation tasks, the objective is to grasp the object in a manner that is configurationally compatible with the downstream manipulation task. Classical TOG methods evaluate the task-oriented grasp quality with task wrench space analysis [22]. However, they share similar limitations with earlier studies in TAG due to the inherent diversity of objects and tasks in open-world environments. Recently, data-driven methods have achieved success in solving TOG problems to some extent. Dang et al.[23] build a task-oriented grasp planner that considers partial object geometry, tactile contacts, and hand kinematic data. Yang et al.[24] expand previous works to enable TOG in object-stacking scenes. To alleviate the burden of data collection, Fang et al.[25] and Qin et al. [26] train a dedicated TOG model for each task with simulated self-supervision. Nevertheless, the aforementioned methods acquire object-task-grasp relationships solely from training data without external knowledge sources and consequently exhibit unsatisfying generalization performance.
More recent works have proposed incorporating external prior knowledge into TOG pipelines. Montesano et al.[3] and Song et al. [4][27] construct semantic knowledge bases (KBs) with pre-defined sets of objects, tasks, actions, and constraints. Similarly, Ardón et al. [28] use Markov Logic Networks (MLN) to build a semantic KB, encoding objects, tasks, visual attributes, and indoor locations of 30 commonly found household objects. To enrich a KB with a broader range of objects and tasks, Murali et al.[5] contribute the largest and most diverse TOG dataset, TaskGrasp, and a resulting KG. More importantly, they propose the state-of-the-art graph-based TOG method, GCNGrasp, to generalize to objects and tasks within the KG. In addition to only incorporating semantic knowledge, a series of works [6, 7] consider both semantic and geometric properties of pre-defined sets of objects and tasks. For example, Detry et al.[6] approach the TOG problem with a geometric grasp model and a CNN-based semantic model. Similarly, in [7], a visual module that estimates object geometric properties is integrated with a probabilistic logic module. Another line of research utilizes affordance knowledge to predict task-oriented grasps. Xu et al. [29] combine affordance segmentation and keypoint detection to guide task-oriented grasp prediction. Liu et al.[30] propose a novel representation of task-oriented grasp contexts with part affordances, part material, object states, and tasks.
While these methods have demonstrated the generalization capability to some extent, they experience significant performance degradation when handling novel objects and tasks out of the pre-defined sets. This phenomenon can be explained by the fact that these methods limit the prior knowledge to a closed-set scope. To learn generalizable TOG skills, we propose to leverage open-ended knowledge from foundation models in this work.
II-C Foundation Models in Robotics
Foundation models, such as LLMs and VLMs, encode open-ended knowledge about the world and have exhibited impressive capabilities on a wide range of visual and linguistic tasks. In robotics, such knowledge is beneficial for physically grounded tasks, such as tabletop manipulation[8, 9], navigation[10, 11], and mobile manipulation[12]. While most of these methods primarily rely on foundation models for task planning and scene understanding, FoundationGrasp grounds the open-ended knowledge to the grasp synthesis process.
Regarding robotic grasping, LERF-TOGO[31] builds a Language Embedded Radiance Fields (LERF) with features distilled from a VLM to produce a 3D relevancy heatmap given an object part name. LAN-Grasp [32] first prompts an LLM to output an object part name appropriate for grasping and uses a VLM to identify the corresponding part in the image. These two methods focus on solving “where to grasp” and ignore “how to grasp”. FoundationGrasp optimizes 6 DoF task-oriented grasp poses to jointly address “where to grasp” and “how to grasp”. GraspCLIP[33] and ATLA[34] extract knowledge from foundation models for manipulation and grasping. However, they are limited to 2D planar grasping. In contrast, FoudationGrasp studies TOG in 3D space.
III Problem Formulation
We consider the problem of learning a task-oriented grasping policy $\pi$ that takes in an input tuple $\gamma=(X_{o},I_{o},L)$ and outputs a task-oriented grasp pose $g^{*}$:
$g^{*}=\pi(\gamma)=\pi(X_{o},I_{o},L)$ (1)
where $X_{o}\in\mathbb{R}^{N\times 3}$ and $I_{o}\in\mathbb{R}^{H\times W\times 3}$ are the (partial) point cloud and the RGB image of the object instance $o$, respectively. $N$ is the number of points. $H$ and $W$ are the height and width of the image. $L$ is a natural language instruction specifying an object class $c$ and a task $t$, such as “pour water with the mug”. $g^{*}$ is a 6 DoF grasp pose of a parallel-jaw gripper represented by $(R,T)$, where $R\in SO(3)$ represents the 3D orientation and $T\in\mathbb{R}^{3}$ represents the 3D translation.
We assume access to a TOG dataset $\mathcal{D}$, consisting of $K_{o}$ object instances, $\mathcal{O}=\{o_{k}\}_{k=1}^{K_{o}}$, with ground truth task-oriented grasp pose annotations. $\mathcal{D}$ spans a set of object classes $\mathcal{C}=\{c_{i}\}_{i=1}^{K_{c}}$ and a set of tasks $\mathcal{T}=\{t_{j}\}_{j=1}^{K_{t}}$, where $K_{c}$ and $K_{t}$ are the numbers of object classes and tasks in $\mathcal{D}$, respectively. We use $\mathcal{D}$ to supervise the training of $\pi$. During inference, $o$, $c$ and $t$ can be novel elements out of $\mathcal{D}$. The key challenge here is to generalize $\pi$ learned on $\mathcal{D}$ to an open set of elements not necessarily trained with.
Mathematically, the goal is to estimate the posterior distribution $P(G|X_{o},I_{o},L)$, where $G$ represents the space of all task-oriented grasp poses. We factorize the estimation process into two steps: (1) task-agnostic grasp sampling $P(G|X_{o},I_{o})$ and (2) task-oriented grasp evaluation $P(S|X_{o},I_{o},L,g)$, where $S$ is the task-compatibility score of each candidate $g\in G$. Finally, the robot executes $g$ with the highest score (i.e., $g^{*}$).
Figure 1: An overview of FoundationGrasp framework: The pipeline consists of (1) data generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. When presented with a language instruction, FoundationGrasp first prompts an LLM to generate semantic and geometric descriptions of the object and the task. A web-based image retrieval module crowdsources images from the Internet. Subsequently, an LLM-based semantic knowledge encoder and a VLM-based geometric knowledge encoder transform them with multi-modal sensory inputs into their feature representations. In the final stage, a Transformer-based task-oriented grasp evaluator with semantic and geometric branches evaluates the task compatibility of each grasp candidate.
IV FoundationGrasp
IV-A Overview
An overview of the proposed FoundationGrasp framework is depicted in Figure 1. The pipeline consists of three stages: (1) data generation, (2) multi-modal feature representation, and (3) task-oriented grasp evaluation. In Section IV-B, FoundationGrasp generates linguistic data (i.e., language descriptions and instructions), visual data (i.e., images), and grasp data (i.e., grasp poses) based on $o$, $c$, and $t$. The second stage described in Section IV-C transforms multi-modal data into their latent space feature representations. Finally, in Section IV-D, a Transformer-based task-oriented grasp evaluator computes the task-compatibility score $S$ of each candidate grasp pose $g$.
IV-B Data Generation
The data generation stage constructs the LaViA-TaskGrasp dataset by augmenting the TaskGrasp dataset with linguistic and visual data. The original TaskGrasp dataset consists of $K_{o}$ object instances, spanning $K_{c}$ object classes and $K_{t}$ tasks. The data generation pipeline consists of four phases: (1) language description generation, (2) language instruction generation, (3) web-based image retrieval, and (4) grasp candidate generation.
Class O2O Description O2T Description
Mug
[sim_geo] “Objects such as teacups, jars, glasses, and
cylindrical vases have similar geometries to a mug.”
[func] “A mug is a typically cylindrical household object with a
handle, used primarily for drinking hot beverages, such
as coffee, tea, or hot chocolate.”
Task T2T Description T2O Description
Sweep
[sim_effect] “Verbs such as clear, clean, brush, wipe, or
dust achieve similar effects to ’sweep an object’.”
[afford] “Brooms, dustpans, vacuum cleaners, and sweepers are
household objects that afford the function of sweeping.”
TABLE I: Examples of semantic descriptions in the LaViA-TaskGrasp dataset
Class O2P Description Task T2P Description
Mug
[part] “A mug is composed of a cylindrical body, a
handle, a circular base, and often features an
outer glaze or design.”
Sweep
[primitive] “Primitive shapes on a household object that
can be used for sweeping include the long, cylindrical
handle and the fan-shaped bundle of bristles on a broom.”
TABLE II: Examples of geometric descriptions in the LaViA-TaskGrasp dataset
Language Description Generation   To extract open-ended knowledge from foundation models, we start by prompting an LLM to generate descriptions for each object class $c_{i}$ and task $t_{j}$ in the TaskGrasp dataset. We divide language description generation into semantic and geometric aspects and carefully design the LLM prompting strategy to capture the relationships between objects, tasks, and grasps.
Semantic descriptions linguistically connect novel objects and tasks with familiar ones. Previous methods [5, 3, 4, 27, 28] primarily model three types of semantic relationships: (1) object-object, (2) task-task, and (3) object-task. We follow a similar spirit to design semantic description prompts. For each object class $c_{i}$, we design two prompt sets: (1) object-to-object (O2O) prompt set, with each prompt querying object classes sharing a similar property with $c_{i}$ (e.g., “Describe what household objects have similar $P_{obj}$ to $c_{i}$:”). Here, $P_{obj}$ represents object properties, such as shape, function, and meta-category; (2) object-to-task (O2T) prompt set, with each prompt querying tasks afforded by $c_{i}$ (e.g., “Describe the common use of household object $c_{i}$:”). Similarly, for each task $t_{j}$, we design (1) task-to-task (T2T) prompt set, with each prompt querying tasks that are semantically or physically relevant to $t_{j}$ (e.g., “Describe what verbs are semantically close to $t_{j}$:”); (2) task-to-object (T2O) prompt set, with each prompt querying object classes that afford $t_{j}$ (e.g., “Describe what household objects can be used to $t_{j}$:”).
Geometric descriptions focus on describing two aspects: (1) the geometric structure of $c_{i}$ and (2) the correlations between geometric structures (form) and $t_{j}$ (function). To address the former, we design an object-to-part (O2P) prompt set, with each prompt querying the geometric structure of $c_{i}$ (e.g., “Describe what parts the household object $c_{i}$ is composed of:”). In parallel, for each task $t_{j}$, we design a task-to-part (T2P) prompt set, with each prompt querying geometric structures that support $t_{j}$ (e.g., “Describe what geometric primitives on a household object can be used to $t_{j}$:”). We regularize the outputs of the LLM to bullets of parts or primitive shapes. This regularization strategy reduces the difficulty of grounding geometric descriptions to 3D representations during task-oriented grasp evaluation.
Each prompt set includes $N_{p}$ prompts, with $N_{p}$ varying for each set. To account for the inherent randomness of generated descriptions, we query the LLM $N_{a}$ times per prompt. Examples of generated semantic and geometric descriptions are showcased in Table I and II. Since a single description only offers a partial “observation” of $c_{i}$ or $t_{j}$, we randomly combine descriptions from each prompt to construct complete semantic description paragraphs, denoted as $L_{c_{i}}^{sem}$ and $L_{t_{j}}^{sem}$, and geometric description paragraphs, denoted as $L_{c_{i}}^{geo}$ and $L_{t_{j}}^{geo}$. Typically, each paragraph has an approximate length of 2-6 sentences. A complete list of prompt sets and more examples of language descriptions can be found in the appendix. During inference, when presented with novel objects and tasks out of the LaViA-TaskGrasp dataset, FoundationGrasp follows the same procedure to generate semantic and geometric descriptions for them. Such an open-ended nature is the primary advantage of FoundationGrasp over previous TOG methods.
Language Instruction Generation  During training, the language goal is constructed from instruction templates derived from [35]. Each template in the initial set, such as “Use the [obj] to [task]”, is populated with an object class $c_{i}$ and a task $t_{j}$. Recognizing that human-expressed language goals exhibit significant diversity, we further enrich the vocabulary and grammatical diversity of the initial template set using the LLM. Specifically, we prompt the LLM to perform template augmentation (e.g., “rewrite the following sentence in a different grammatical format:”), which leads to an augmented set of instruction templates. The augmentation process generates some unconventional formulations (e.g., “ensure you have a [task]-friendly grip on the [obj]”). While not commonly used in daily communication, we keep them to enhance the robustness to novel instructions. A complete list of instruction templates is available in the appendix.
Web-Based Image Retrieval  Previous works in 3D scene representation have demonstrated that visual-language features distilled from 2D foundation models facilitate geometric understanding of 3D world [36, 37]. Following the generation of geometric descriptions, we utilize the Internet as a large-scale database to retrieve visual data. Given that each instance in the TaskGrasp dataset is captured from only 2-3 viewpoints, we augment the dataset by crowdsourcing images from the Internet. Specifically, for each object instance $o_{k}$ in the dataset, we first query Google Cloud Vision API with its original images in the dataset, retrieving candidate images with respective confidence scores. Each candidate image contains an object instance sharing similar visual characteristics and geometric structures. Subsequently, we employ the Segment Anything Model (SAM) [38] for background removal. To guarantee the view diversity, we utilize CLIP [39] visual encoder to extract the embedding vector of each candidate image and compute the cosine similarities between them. Those with high similarities are excluded. Finally, a set of $N_{I}$ images (original + retrieved images) for $o_{k}$, denoted as $I_{o_{k}}$, is selected based on confidence scores. Examples of retrieved images are available in the appendix. Note that the goal here is not to perform accurate 3D reconstruction of $o_{k}$. Thus, the instances in the retrieved images do not need to match $o_{k}$ precisely. This image retrieval strategy aims to provide visual information for constructing geometric priors of objects and tasks in the later stage. During inference, the robot utilizes $I_{o}$ captured by the onboard camera.
Grasp Candidate Generation  Each object instance $o_{k}$ in the TaskGrasp dataset is annotated with 25 candidate grasp poses and task labels. These annotations supervise the training of $\pi$. During inference, we first apply Contact-GraspNet [21] to generate $N_{g}^{+}$ collision-free grasp candidates on $X_{o}$. $N_{g}^{+}$ varies depending on the specific scene. Then, a subset of $N_{g}$ grasps ($N_{g}\leq N_{g}^{+}$) is selected using farthest point sampling (FPS) to construct the final set of candidate grasps. The selection of $N_{g}$ is a trade-off between grasp diversity and density.
Figure 2: A grasp $g$ is represented with six control points $X_{g}$ on the gripper model in the object reference frame. The origin of the object frame is $\overline{X}$, the center of mass of the object point cloud $X_{o}$.
IV-C Multi-Modal Feature Representation
To incorporate open-ended knowledge from the previous stage into the FoundationGrasp framework, we transform them with multi-modal sensory inputs into their latent space feature representations. Three encoders are introduced: an object and grasp encoder, an LLM-based semantic knowledge encoder, and a VLM-based geometric knowledge encoder.
Object and Grasp Encoder  To model the relative spatial relationship between the gripper and the object, the first step involves transforming $g\in SE(3)$ and $X_{o}\in\mathbb{R}^{N\times 3}$ into the same space. This is achieved by approximating $g$ with six control points[19], represented as $X_{g}\in\mathbb{R}^{6\times 3}$ in the object frame (shown in Figure 2). Then, $X_{o}$ and $X_{g}$ are jointly embedded. Specifically, $X_{o}$ and $X_{g}$ are first concatenated to form $X_{o+g}\in\mathbb{R}^{(N+6)\times 3}$. A binary feature vector is appended to indicate whether a point belongs to $X_{o}$ or $X_{g}$. Subsequently, a PointNet++[40] like encoder (denoted as PN++) extracts localized features from $X_{o+g}$, which associates features to a reduced number of $N^{-}$ points ($N^{-}<N$):
$X_{o+g}$ $=\textnormal{Concat}([X_{o},X_{g}],\textnormal{dim}=0)$ (2)
[…]
where point embeddings $F_{X}\in\mathbb{R}^{N^{-}\times 1024}$. We choose to maintain a degree of locality in the latent space, as opposed to relying on a single global embedding, to preserve the localized geometric information. Such information is necessary for understanding the relationship between geometric structures, objects, and tasks, as discussed in Section IV-B. The point encoder consists of three set abstraction (SA) layers with multi-scale grouping (MSG). A Multi-Layer Perceptron (MLP) is then applied for dimension reduction to output $\tilde{F}_{X}\in\mathbb{R}^{N^{-}\times 256}$.
Semantic Knowledge Encoder  To linguistically connect novel and familiar objects and tasks, FoundationGrasp needs to interpret a broad spectrum of linguistic concepts present in semantic descriptions. Consider, for instance, a T2O description of task “scoop”: “Objects such as spatulas, ladles, and spoons are kitchen utensils that support the function of scoop.” In this context, FoundationGrasp needs to identify object classes (e.g., spatula, ladle, spoon), understand their categorical taxonomy (e.g., utensil), and recognize associated tasks (e.g., scoop). Training a dedicated language encoder from scratch would take substantial training effort and, meanwhile, achieve limited generalization to novel linguistic concepts unseen during training. As an alternative, we utilize a large pre-trained BERT[41] to encode $L_{c_{i}}^{sem}$ and $L_{t_{j}}^{sem}$. Other large pre-trained language models would also suffice. These LLMs are pre-trained on a large corpus of text data so that they can interpret and generalize to novel linguistic concepts without any fine-tuning. Specifically, the pre-trained BERT processes a paragraph from $L_{c_{i}}^{sem}$ to output $F_{c_{i}}^{sem}\in\mathbb{R}^{T_{c}\times 768}$, and a paragraph from $L_{t_{j}}^{sem}$ to output $F_{t_{j}}^{sem}\in\mathbb{R}^{T_{t}\times 768}$. Likewise, we transform a language instruction $L$ into $F_{L}\in\mathbb{R}^{T_{L}\times 768}$. $T_{c}$, $T_{t}$, and $T_{L}$ denote the maximum lengths (with zero-padding) of $L_{c}^{sem}$, $L_{t}^{sem}$, and $L$, respectively. We further project language sequence embeddings to a lower dimension and obtain $\tilde{F}_{c_{i}}^{sem}\in\mathbb{R}^{T_{c}\times 512}$, $\tilde{F}_{t_{j}}^{sem}\in\mathbb{R}^{T_{t}\times 512}$, and $\tilde{F}_{L}\in\mathbb{R}^{T_{L}\times 512}$.
Geometric Knowledge Encoder  Geometric knowledge is twofold, including images and geometric descriptions. Images visually capture the geometric structure of an object, while geometric descriptions provide a linguistic interpretation of the object’s geometry. Similar to previous works [36, 37], we extract multi-modal geometric features with CLIP, which is pre-trained on a large number of geometric concepts (e.g., primitives, parts, and objects) in the form of image-text pairs. During each training loop, we first extract intermediate CLIP visual features of a random image from $I_{o_{k}}$ to obtain $F_{I_{o_{k}}}\in\mathbb{R}^{h\times w\times 1024}$. Paragraphs from $L_{c_{i}}^{geo}$ and $L_{t_{j}}^{geo}$ are then processed by the CLIP language encoder to obtain $F_{c_{i}}^{geo}\in\mathbb{R}^{1024}$ and $F_{t_{j}}^{geo}\in\mathbb{R}^{1024}$, respectively. Then, flattening and dimension reduction are performed to output $\tilde{F}_{I_{o_{k}}}\in\mathbb{R}^{hw\times 128}$, $\tilde{F}_{c_{i}}^{geo}\in\mathbb{R}^{128}$, and $\tilde{F}_{t_{j}}^{geo}\in\mathbb{R}^{128}$.
Figure 3: Task-oriented grasp evaluator is a customized Transformer consisting of a geometric branch $TGE_{geo}$ (left) and a semantic branch $TGE_{sem}$ (right).
IV-D Task-Oriented Grasp Evaluation
The final stage evaluates the task compatibility of $g$, leveraging the semantic and geometric knowledge embedded in multi-modal features. To this end, we propose a Transformer-based task-oriented grasp evaluator with semantic and geometric branches. For simplicity, we drop the subscripts of multi-modal features in the following discussions. Formally, task-oriented grasp evaluation can be formulated as:
$S=\textnormal{TGE}(\tilde{F}_{X},\tilde{F}_{I},\tilde{F}_{c}^{geo},\tilde{F}_{t}^{geo},\tilde{F}_{L},\tilde{F}_{c}^{sem},\tilde{F}_{t}^{sem})$ (4)
where TGE represents the task-oriented grasp evaluator, and $S$ is the task-compatibility score of $g$. The first four inputs are processed by the geometric branch $\textnormal{TGE}_{\textnormal{geo}}$, and the latter three are processed by the semantic branch $\textnormal{TGE}_{\textnormal{sem}}$.
Semantic Branch  $\textnormal{TGE}_{\textnormal{sem}}$ aims to linguistically connect novel objects and tasks with familiar ones using semantic descriptions. The problem is analogous to machine translation in natural language processing (NLP), which aims to transform language sequences from the source to the target domain. In our context, the robot initially does not understand novel elements in $L$. $\textnormal{TGE}_{\textnormal{sem}}$ translates them to familiar ones using semantic descriptions. Inspired by this similarity, we design $\textnormal{TGE}_{\textnormal{sem}}$ based on the Transformer decoder, an effective architecture for handling sequence-to-sequence problems.
The translation process is divided into two phases: object knowledge incorporation and task knowledge incorporation, with each phase being addressed by a Transformer decoder layer. The architecture of $\textnormal{TGE}_{\textnormal{sem}}$ is presented in Figure 3 (right), which consists of $\phi^{sem}_{c}$ and $\phi^{sem}_{t}$. Each layer incorporates semantic description embeddings $\tilde{F}^{sem}$ (i.e., $\tilde{F}_{c}^{sem}$ or $\tilde{F}_{t}^{sem}$) into $\tilde{F}_{L}$:
$\phi^{sem}(\tilde{F}^{sem},\tilde{F}_{L})\rightarrow\tilde{F}_{L}^{sem}\in\mathbb{R}^{T_{L}\times 512}$ (5)
where $\tilde{F}_{L}^{sem}$ denotes language instruction embeddings contextualized with semantic knowledge. The computational procedure of the two layers is similar, which can be represented as:
$\tilde{F}_{L}^{sem}$ $=\text{LN}(\tilde{F}_{L}+\text{MHA}(\tilde{F}_{L},\tilde{F}^{sem}))$ (6a)
$\tilde{F}_{L}^{sem}$ $=\text{LN}(\tilde{F}_{L}^{sem}+\text{FFN}(\tilde{F}_{L}^{sem})$ (6b)
where LN, MHA, and FFN denote layer normalization, multi-head attention, and feedforward networks, respectively. MHA consists of $h_{sem}$ heads. Each head reconstructs $\tilde{F}_{L}$ with elements in $\tilde{F}^{sem}$ weighted by their normalized correspondence:
$\text{head}_{i}$ $=\text{Attention}(Q_{L},K_{sem},V_{sem})$ (7a)
$=\text{Softmax}(\frac{Q_{L}K_{sem}^{T}}{\sqrt{d_{k}}})V_{sem}$ (7b)
where $n=1,2,...,h_{sem}$. $Q_{L}$, $K_{sem}$, and $V_{sem}$ can be computed as:
$Q_{L}=\tilde{F}_{L}W_{n}^{Q},K_{sem}=\tilde{F}^{sem}W_{n}^{K},V_{sem}=\tilde{F}^{sem}W_{n}^{V}$ (8)
where $W_{n}^{Q}\in\mathbb{R}^{512\times d_{k}}$, $W_{n}^{K}\in\mathbb{R}^{512\times d_{k}}$, $W_{n}^{V}\in\mathbb{R}^{512\times d_{v}}$, and $d_{k}=d_{v}$. During training, semantic descriptions are associated with corresponding object and task tokens, while irrelevant information is disregarded. Finally, $\tilde{F}_{L}^{sem}$ is mean pooled to obtain the language instruction embedding $\overline{F}_{L}^{sem}\in\mathbb{R}^{128}$ contextualized with semantic knowledge.
Geometric Branch  To facilitate the geometric understanding of objects and tasks, $\textnormal{TGE}_{\textnormal{geo}}$ fuses multi-modal geometric features distilled from 2D VLM to 3D. Specifically, to match the spatial dimension of $\tilde{F}_{I}$, the geometric description embeddings are first repeated at each spatial location to obtain $\tilde{F}_{c}^{geo}\in\mathbb{R}^{hw\times 128}$ and $\tilde{F}_{t}^{geo}\in\mathbb{R}^{hw\times 128}$. Then, multi-modal geometric features are constructed as below:
[…]
where $\tilde{F}_{I}^{geo}\in\mathbb{R}^{hw\times 256}$ ($\tilde{F}_{I,c}^{geo}$ or $\tilde{F}_{I,t}^{geo}$ ). The architecture of $\textnormal{TGE}_{\textnormal{geo}}$ is depicted in Figure 3 (left). Similar to the semantic branch, the geometric branch is divided into two object and task knowledge incorporation, which correspond to $\phi_{c}^{geo}$ and $\phi_{t}^{geo}$ in $\textnormal{TGE}_{\textnormal{geo}}$, respectively. Each layer incorporates geometric features $\tilde{F}^{geo}_{I}$ into $\tilde{F}_{X}$:
$\phi^{geo}(\tilde{F}^{geo}_{I},\tilde{F}_{X})\rightarrow\tilde{F}_{X}^{geo}\in\mathbb{R}^{N^{-}\times 256}$ (10)
The two layers share a similar design, where geometric features are fused with a cross-attention layer and rectified by a self-attention layer:
$\tilde{F}_{X}^{geo}$ $=\textnormal{LN}(\tilde{F}_{X}+\textnormal{MHA}(\tilde{F}_{X},\tilde{F}_{I}^{geo}))$ (11a)
$\tilde{F}_{X}^{geo}$ $=\textnormal{LN}(\tilde{F}_{X}^{geo}+\textnormal{MHA}(\tilde{F}_{X}^{geo}))$ (11b)
where MHA consists of $h_{geo}$ heads. The computational procedure of cross-attention fusion is similar to Equation 6. The self-attention rectification operates similarly to cross-attention fusion, with the distinction that query, key, and value tensors are all projections of $\tilde{F}_{X}^{geo}$. The design choice, where a rectification layer follows a fusion layer, is based on the consideration that a self-attention layer performs a permutation-invariant transformation to rectify geometric features that are not properly aligned[42]. Finally, $\tilde{F}_{X}^{geo}$ is processed by an SA layer with global pooling to output the point embedding $\overline{F}_{X}^{geo}\in\mathbb{R}^{300}$ contextualized with geometric knowledge.
Finally, the language instruction embedding $\overline{F}_{L}^{sem}$ is concatenated with the point embedding $\overline{F}_{X}^{geo}$ to compute $S$, which can be represented as:
[…]
where $\sigma$ denotes the sigmoid activation function.
Loss Function  The binary cross-entroy loss between $S$ and the ground truth label $S_{gt}$ is computed as below:
$\mathcal{L}$ $=\textnormal{BCE}(S,S_{gt})$ (13a)
$=-\frac{1}{M}\sum_{m=1}^{M}S_{gt,m}\cdot\text{log}(S_{m})+(1-S_{gt,m})\cdot\text{log}(1-S_{m})$ (13b)
where $M$ is the number of samples, and $S_{gt,m}$ is set to one if the $m_{th}$ grasp pose is compatible with $t$ and zero otherwise.
V Experimental Setup
In this section, we evaluate the performance of FoundationGrasp against existing baselines under three different held-out settings, including held-out instance setting, held-out (object) class setting, and held-out task setting. We divide the experiments into two parts: perception experiments and real-robot experiments. While the former evaluates the performance on the LaViA-TaskGrasp dataset, the latter is conducted on a real robot. Lastly, implementation details for model design and training are provided.
V-A Perception Experiments
Baselines  We compare FoundationGrasp to the following baselines that are representative of existing grasp synthesis methods:
  • TAG represents task-agnostic grasping methods [17, 18, 19, 20, 21] that focus on satisfying the stability constraint while ignoring the task-compatibility constraint. Specifically, we use Contact-GraspNet[21] as our baseline.
  • SGN is an adapted version of Semantic Grasp Network proposed in [30]. It learns object-task-grasp relationships solely from training data without incorporating external knowledge.
  • SGN+we is an enhanced version of SGN. It uses ConceptNet[43] pre-trained word embeddings as prior knowledge to initialize object classes and tasks.
  • GCNGrasp [5] is the state-of-the-art graph-based TOG method. It incorporates a semantic KG connecting a pre-defined set of objects and tasks. In our implementation, given a novel object class or task out of the graph, we connect the node to its nearest neighbor in the graph by measuring the cosine similarity of their ConceptNet pre-trained word embeddings.
We also report the performance of GraspGPT to highlight the technical extension from the preliminary version.
Figure 4: Test objects in real-robot experiments cover commonly used accessories, kitchen utensils, and mechanic tools.
Dataset  FoundationGrasp and baselines are evaluated on the contributed LaViA-TaskGrasp dataset, which includes 56 tasks, 75 object classes, and 191 object instances. Each object instance is a point cloud of a real household object reconstructed from multiple viewpoints. 25 candidate grasp poses with binary labels (1 for task-compatible, 0 for task-incompatible) are annotated on each object instance. For each held-out setting, $\mathcal{K}$-fold cross validation ($\mathcal{K}=4$) is performed. In each fold, 25% of instances/classes/tasks are used for testing, and the remaining 75% are used for training and validation. In terms of language data, the LaViA-TaskGrasp dataset contains 40 geometric descriptions, 80 semantic descriptions for each object class, and 40 geometric descriptions, 50 semantic descriptions for each task. We randomly combine these descriptions to generate 750 object semantic/geometric description paragraphs and 560 task semantic/geometric description paragraphs. LaViA-TaskGrasp dataset also includes 53 language instruction templates, resulting in a diverse set of over two million possible language instructions. In terms of vision data, we prepare $N_{I}=24$ images, as in [42], for $o_{k}$ following the procedure described in Section IV-B.
Metrics  According to the convention in previous works[5, 14], we compute the Average Precision (AP) scores for each object instance, object class, and task. These scores are then averaged over all instances, classes, and tasks to obtain instance mAP (mean AP), class mAP, and task mAP.
Method Held-out Instance Performance Held-out Class Performance Held-out Task Performance
Instance mAP Class mAP Task mAP Instance mAP Class mAP Task mAP Instance mAP Class mAP Task mAP
TAG 59.75 60.28 54.76 59.32 58.73 52.72 59.06 58.24 52.37
SGN 78.51 75.08 68.80 74.20 72.95 62.55 75.17 71.59 63.35
SGN+we 79.74 77.91 74.36 77.21 75.51 63.73 78.06 74.49 70.55
GCNGrasp 83.24 82.02 73.71 72.92 72.45 67.58 57.48 47.21 33.17
GraspGPT 80.25 77.94 76.10 79.70 77.88 72.84 79.32 76.90 72.34
FoundationGrasp 93.17 91.55 89.39 88.74 86.09 84.33 89.01 87.36 82.70
TABLE III: Quantitative results of perception experiments
V-B Real-Robot Experiments
Task-Oriented Grasping Experiments  The experiment platform comprises a 7 DoF Kinova Gen3 robotic arm equipped with a Robotiq 2F-85 parallel jaw gripper and an Intel RealSense D435 RGB-D camera with eye-in-hand calibration. Shown in Figure 4, we collect test objects from the YCB dataset and our laboratory, covering commonly used accessories, kitchen utensils, and mechanic tools. During inference, SAM [38] is first applied to extract the test object from the point cloud captured from a single view. Contact-GraspNet [21] then samples $N_{g}^{+}$ grasp candidates for evaluation. Compared to the multi-view setting in perception experiments, the single-view setup is used in real-robot experiments since a robot, in most cases, only has a partial observation of the object.
Task-Oriented Manipulation Experiments  To support the hypothesis that FoundationGrasp facilitates downstream tool manipulation, we combine FoundationGrasp with pre-programmed manipulation skills and conduct task-oriented manipulation experiments. Presented in Table V, we conduct manipulation experiments on nine household tasks selected from the household chores list in [44]. Since manipulation policy learning is not the focus of this paper, we design rule-based heuristics to determine the operating direction of the tool object and compute the effect point on the receptacle object. Manipulation trajectories are generated by merging sequential motion primitives.
Metrics  The task-oriented grasping pipeline is divided into three stages: perception, planning, and action. We report the statistics of each stage separately, providing details for failure analysis. A trial is considered successful if the test object is grasped by the robot in a task-oriented manner given a language instruction. For task-oriented manipulation experiments, we record the success of task-oriented grasping and manipulation, respectively.
V-C Implementation Details
Training and inference are conducted on a desktop PC with a single NVIDIA RTX 3090 GPU. All the baselines are re-trained on the LaViA-TaskGrasp dataset for fair comparison. FoundationGrasp is implemented using Pytorch 1.11.0 with CUDA 11.3. We use an Adam [45] optimizer with a weight decay of 0.0001. The learning rate is set to 0.0001 initially and adjusted with a LambdaLR scheduler as in [5, 14]. To increase the robustness to unseen object poses, input point clouds are augmented with random scaling, rotation, jitter, and dropout. Each point cloud is then downsampled to 4096 points and fed into FoundationGrasp. The model is trained for 50 epochs with a batch size of 32.
For the choice of foundation models, we use OpenAI GPT-4 to generate language descriptions and OpenAI CLIP implemented by OpenCLIP as the VLM. For the large-pretrained language encoder, we choose the Google pre-trained BERT-Base model provided by Hugging Face. All the foundation models are frozen during training. More details on model training and design can be found in the appendix.
Figure 5: Qualitative results of perception experiments. GT denotes ground truth task-oriented grasp poses. FoundationGrasp, GCNGrasp, and GraspGPT are evaluated under three held-out settings. Top-5 predictions of each method are visualized. Grasp poses are colored by their confidence scores (green is higher).
Figure 6: FoundationGrasp exhibits commonsense reasoning capability. It rejects all grasp candidates under conditions where the object is incompatible with the intended task, or there is a discrepancy between the target and actual object classes.
VI Experiments
VI-A Results of Perception Experiments
The quantitative results of perception experiments are reported in Table III. For three held-out settings, TAG establishes a lower performance bound for all methods, achieving approximate mAPs of 50-60%. The seemingly high mAPs by random guessing is because the distribution of positive and negative samples in the LaViA-TaskGrasp dataset is about even. In the case where the dataset is annotated with more negative samples, mAPs are expected to be much lower. SGN surpasses TAG’s performance by over 10% since it considers the task constraints on object grasping. Moreover, ConceptNet word embeddings capture the semantic similarity between relevant concepts in the embedding space. Consequently, SGN+we achieves a marginal performance boost compared to SGN.
GCNGrasp performs decently well, with mAP scores of 83.24%, 82.02%, and 73.71%, when dealing with intra-class variances. It can be attributed to the fact that the semantic KG pre-defines relations between all classes and tasks in the dataset. Nonetheless, GCNGrasp experiences significant performance drops when generalizing to novel classes and tasks out of KG due to its inherent closed-world nature. Such a limitation becomes especially evident in the held-out task setting, where GCNGrasp even falls behind TAG by 1.58%, 11.03%, and 19.20%, and SGN by 17.69%, 24.38%, and 30.18% across three evaluation metrics. It also suggests that ConceptNet word embeddings are more suitable for capturing linguistic relations between objects than tasks.
FoundationGrasp and GraspGPT outperform TAG, SGN, and GCNGrasp in both the held-out class and task settings, demonstrating the superiority of the open-ended design. Due to the incorporation of geometric knowledge, FoundationGrasp further improves the performance of GraspGPT by approximately 10%. Overall, FoundationGrasp performs best in all held-out settings. Regarding running time, it takes less than two seconds for FoundationGrasp to evaluate 25 grasp candidates on a single RTX 3090. A qualitative comparison between all methods is illustrated in Figure 5. Notably, even though baselines may assign high confidence scores to their predictions, they often fail to accurately localize the task-oriented regions. FoundationGrasp also exhibits commonsense reasoning capability to some degree. As is depicted in Figure 6, FoundationGrasp rejects all grasp candidates (i.e., assign low confidence scores) under conditions where the object is incompatible with the intended task, or there is a discrepancy between the target and actual object classes.
VI-B Results of Real-Robot Experiments
Task-Oriented Grasping Experiments  The result of task-oriented grasping experiments is summarized in Table IV. 100 trials are conducted for each held-out setting, with 10 per instance, class, or task. Despite capturing object point clouds from a single viewpoint, FoundationGrasp consistently achieves a perception success rate exceeding 80% across all settings. Similar to earlier observations in perception experiments, GCNGrasp suffers when generalizing to novel classes and tasks, while GraspGPT and FoundationGrasp perform consistently better. The improvement observed with FoundationGrasp over GraspGPT indicates the necessity of incorporating both semantic and geometric knowledge. FoundationGrasp experiences a significant performance drop from the perception to the action stage. Three typical failure modes are (1) unstable grasp candidates generated by the task-agnostic grasp sampler, (2) motion planning failures, and (3) false negative evaluations by FoundationGrasp (e.g., rejecting all valid candidates). Concerning the third failure mode, the training stage of FoundationGrasp can be conceptualized as constructing an implicit knowledge base in the latent space with a set of objects and tasks. During the inference stage, FoundationGrasp aims to build connections between existing elements in the knowledge base and novel ones outside. However, if novel elements are substantially divergent from the knowledge base, FoundationGrasp tends to make false negative predictions. As illustrated in Figure 8, in the case of instructing the robot to screw with a power drill, object class power drill shares minimal semantic or geometric similarity to existing elements in the knowledge base, leading to the rejection of all valid candidates. Nevertheless, FoundationGrasp achieves the highest grasping success rate, averaging 72.67% over three held-out settings.
Figure 7: Qualitative results of task-oriented manipulation experiments. The robot predicts task-oriented grasps with FoundationGrasp and generates manipulation trajectories by merging sequential motion primitives.
Method Held-out Instance Performance Success Held-out Class Performance Success Held-out Task Performance Success
Perception Planning Action Perception Planning Action Perception Planning Action
GCNGrasp 81/100 70/100 65/100 65.00% 50/100 44/100 37/100 37.00% 50/100 42/100 34/100 34.00%
GraspGPT 83/100 69/100 64/100 64.00% 80/100 69/100 65/100 65.00% 75/100 66/100 59/100 59.00%
FoundationGrasp 89/100 82/100 74/100 74.00% 81/100 77/100 71/100 71.00% 83/100 78/100 73/100 73.00%
TABLE IV: Quantitative results of task-oriented grasping experiments
Figure 8: Since the novel object class “Power Drill” shares minimal similarity to familiar elements in the implicit knowledge base, FoundationGrasp makes false positive predictions by rejecting all valid candidates. The colors of novel elements denote the level of similarity to the “Mug”.
Task-Oriented Manipulation Experiments  This evaluation aims to prove that task-oriented grasping facilitates downstream manipulation tasks. Therefore, baselines are not evaluated here. As reported in Table V, FoundationGrasp achieves a high grasping success rate of approximately 80% on nine manipulation tasks. Specifically, in the Bottle Opening task, since the bottle cap is flat and thin, the robot attempts to grasp at the cap-neck intersection in some cases, resulting in 5 failures out of 10 trials. In terms of manipulation, despite simplifying the manipulation motions into motion primitives guided by rule-based heuristics, FoundationGrasp performs reasonably well, with an average success rate of 66.67%. Two examples of task-oriented manipulation experiments are presented in Figure 7. The performance gap between the grasping and manipulation stage is primarily due to two factors: (1) limited ability to model the relative pose between the tool and the receptacle objects; (2) absence of force feedback in tasks requiring precise contact control, such as Mug Cleaning, Ketchup Squeezing. Future work will consider extending FoundationGrasp to jointly optimize grasp poses and manipulation trajectories and fusing force modality.
Household Task Household Object Success
Grasping Manipulation
Dish Washing Plate, Scrubber 8/10 7/10
Bean Scooping Spoon, Pot 9/10 9/10
Nail Pounding Nail, Hammer 9/10 6/10
Knife Handover Knife 8/10 8/10
Bean Pouring Mug, Bowl 8/10 6/10
Bottle Opening Bottle 5/10 5/10
Ingredients Mixing Spoon, Pot 9/10 7/10
Mug Cleaning Mug, Sponge Brush 7/10 6/10
Ketchup Squeezing Squeeze Bottle, Plate 8/10 6/10
Total 78.89% 66.67%
TABLE V: Quantitative results of task-oriented manipulation experiments
VI-C Ablation Study
To gain further insights into the design choice of FoundationGrasp, we conduct two sets of ablation studies to answer the following questions:
  • How the performance of FoundationGrasp is affected by semantic and geometric branches?
  • How do object and task knowledge help the generalization to novel elements out of the training set?
Ablation on Semantic and Geometric Branches  In this study, we compare FoundationGrasp to three ablated versions: (1) with semantic branch only (i.e., w/ Sem); (2) with geometric branch only (i.e., w/ Geo); (3) vanilla model (i.e., w/o Sem & Geo). The result is presented in Figure 9 (a). Here, we only show mAPs of corresponding held-out settings. For all held-out settings, the full model achieves the best performance, while the vanilla model performs the worst. This demonstrates the effectiveness of semantic-geometric branch architecture. Specifically, for the held-out instance setting, where objects only have limited variances in terms of geometric and visual features, geometric knowledge is essential for unifying such variances. The relations between tasks are easier to capture when tasks are abstracted as linguistic concepts. Therefore, semantic knowledge contributes more to relating different tasks in the held-out task setting. Semantic and geometric branches play an equally important role when generalizing to novel object classes.
Ablation on Object and Task Knowledge  In this study, we compare FoundationGrasp to three ablated versions: (1) with object knowledge only (i.e., w/ Obj); (2) with task knowledge only (i.e., w/ Task); (3) vanilla model (w/o Obj & Task). The result is presented in Figure 9 (b). It is clear from held-out class and task settings that object and task knowledge are essential for generalizing to novel object classes and tasks, respectively. This aligns with the intuition. The full model only outperforms w/ Obj by 1.89% in the held-out class setting. For the held-out instance setting, w/ Obj and w/ Task achieve similar performances (88.47% and 86.98%). Overall, incorporating both object and task knowledge outperforms all ablated versions.
Figure 9: Ablation study: (a) ablation on semantic and geometric branches; (b) ablation on object and task knowledge.
VI-D User Study
To deploy FoundationGrasp on household robots operating in everyday life, it is also necessary to study the practical applicability from the user perspective (e.g., whether the behavior of FoundationGrasp aligns with user needs, preferences, safety, and comfort). Therefore, we conduct a user study with 10 participants on 10 test objects selected from the real-robot experiments benchmark. Once all objects are tested, participants are invited to fill out a questionnaire with Likert scale. Participants are also encouraged to share their experiences or describe any problems they encounter during the user study.
Table VI reports the objective metrics of the user study. On average, the robot can successfully grasp each object with less than two attempts. Figure 10 presents the subjective responses of participants to the questionnaire. Average ratings in all dimensions are above 3 (i.e., Agree or Strongly Agree). Most participants agree that the robot is aware of the task-oriented effect and is able to grasp objects in a task-oriented manner. Some comment that FoundationGrasp exhibits high repeatability (i.e., generates grasp poses around the same point) given the same object and task with different expressions. Participants also note that FoundationGrasp is robust to arbitrary object poses. Such a desirable property is due to the data augmentation performed during training. Regarding safety, one participant expresses concerns about potential collisions with users, resulting in a disagreement on the third statement. Overall, participants suggest the system could be improved in the following ways: (1) The system should consider the state of the object. For example, the generated grasps should be far away from the body of the pot if the pot is hot. (2) The system should be able to understand language instructions directly specifying the part (e.g., “grasp the handle of the pot”).
Hammer Brush Scissors Cereal Box Saucepan Mustard Bottle Spatula Bottle Brush Spoon Screwdriver
Number of Attempts 1.3$\pm$0.15 1.2$\pm$0.20 1.4$\pm$0.22 1.5$\pm$0.27 1.3$\pm$0.15 1.7$\pm$0.26 1.1$\pm$0.10 1.0 1.5$\pm$0.40 1.4$\pm$0.40
Success Rate 70% 90% 70% 70% 70% 50% 90% 100% 80% 80%
TABLE VI: Objective evaluation of user study
VII Discussions and Limitations
Despite compelling results, we observe typical failure modes or limitations during experiments. This section discusses these limitations from data, algorithm, and hardware perspectives and provides insights into future work.
Grasp Data Acquisition  While FoundationGrasp achieves superior generalization performance compared to previous works, the data collection phase still requires labor-intensive manual annotation. Besides, some annotations do not align well with how humans perform task-oriented grasping, which may further hurt the performance. This point can be reflected by the neutral attitude of several participants towards the statement “The robot was able to perform grasping in human-like behavior” in the user study. One potential solution is to learn from human demonstrations. Previous works have proved that human demonstrations (e.g., images or videos) encode constraints and semantics of human behaviors for robot skill learning. Learning with general-purpose simulators of the physical world is also an exciting direction. Recent advances in AIGC (e.g., Sora[46]) open up the potential for generating synthetic demonstration data without a human in the loop.
Sampling-and-Evaluation Strategy  FoundationGrasp follows a two-stage, sampling-and-evaluation pipeline. Using the off-the-shelf task-agnostic grasp sampler reduces the complexity of constructing a TOG pipeline. However, such a design also leads to two problems. First, we assume all grasp candidates from the sampling phase are stable and treat them equally in the evaluation phase. However, we observe that FoundationGrasp tends to bias marginal grasp poses, for example, near the end of the handle. Although functionally correct, these grasp poses are not configurationally robust to perturbations, such as unexpected contact or calibration error. This would lead to potential grasping failure. Second, the first stage samples grasps uniformly over the point cloud. However, robots would only interact with specific regions of an object. Exhaustively evaluating every candidate grasp significantly increases the computational cost of FoundationGrasp and hinders real-time prediction. Future work plans to merge task-agnostic sampling with task-oriented evaluation into a unified framework.
Robot Gripper Generalization  Regarding hardware setup, we currently only consider the parallel jaw gripper. Therefore, FoundationGrasp cannot transfer to other types of grippers, such as three-finger grippers and five-finger dexterous hands. Although such a methodology can be generally applied to training any gripper-dependent TOG model, it is still far from a universal TOG method in terms of both software and hardware. Ideally, the algorithm would take both object geometry and gripper attributes as inputs and output the grasp poses or contact points. The long-term goal is to develop a universal TOG method that can be readily applied to any household robot with any type of gripper.
Figure 10: Participants’ agreement with each statement in the user study questionnaire. Horizontal axis denotes levels of agreement w.r.t. each statement, where Strongly Agree-(5), Agree-(2.5), Neither Agree or Disagree-(0), Disagree-(-2.5), and Strongly Disagree-(5).
VIII Conclusion
In this paper, we present FoundationGrasp, a foundation model-based TOG framework. Compared to previous works, which limit the prior knowledge to a closed-set scope, FoundationGrasp leverages the open-ended knowledge from foundation models to learn generalizable TOG skills. Evaluation on the LaViA-TaskGrasp dataset demonstrates the superiority of FoundationGrasp over existing TOG methods when generalizing to novel object instances, object classes, and tasks. Furthermore, the effectiveness of FoundationGrasp is validated in task-oriented grasping and manipulation experiments on a 7 DoF robotic arm. Lastly, limitations and potential solutions are thoroughly discussed to provide venues for future work.